# How Temporal Unrolling Supports Neural Physics Simulators

### v1

@@ 4.1 Kuramoto-Sivashinsky
This equation is a fourth-order chaotic PDE. The domain size, which leads to more chaotic behavior and shorter Lyapunov times for larger values [15], was varied across training and test data sets. The ground truth solver uses an exponential RK2 integrator [12]. In order to create a challenging learning scenario, the base solver for correction resorts to a first-order version that diverges within 14 steps on average. We base most of our empirical tests on this Kuramoto-Sivashinsky (KS) case since it combines challenging learning tasks with a small computational footprint. The KS equation is a fourth-order stiff PDE governed by
@@
with simulation state $u$, time $t$, and space $x$ living in the domain of size $\mathcal{X}$. The dynamics exhibit a highly chaotic behavior. The domain was discretized with 48 grid points, and timesteps were set to $1$. The physical domain length $\mathcal{X}$ is the critical parameter in the KS equation. The training dataset was computed for a range of $\mathcal{X}=2\pi*[5.6,6.4,7.2,8,8.8,9.6,10.4]$. A sequence of $5000$ steps was computed for each domain length. The numerical solves and network training were performed in PyTorch [45].
@@
We use two architectures for learning tasks with the KS case: a graph-based (GCN) and a convolutional network (CNN). Both networks follow a ResNet-like structure [23], have an additive skip connection from input to output, and are parameterized in terms of their number of features per message-passing step or convolutional layer, respectively. They receive 2 channels as input, a normalized domain size $\mathcal{X}$ and $u$, and produce a single channel, the updated $u$, as output. The GCN follows the hierarchical structure of directional Edge-Conv graph nets [63, 31] where each Edge-Conv block is comprised of two message-passing layers with an additive skip connection. The message-passing concatenates node, edge, and direction features, which are fed to an MLP that returns output node features. At a high level, the GCN can be seen as the message-passing equivalent of a ResNet. These message-passing layers in the GCN and convolutional layers in the CNN are scaled in terms of whole ResBlocks, i.e., two layers with a skip connection and leaky ReLU activation. The networks feature additional linear encoding and decoding layers for input and output. The architectures for GCN and CNN were chosen such that the parameter count matches, as listed in table 2.
@@
Hence, this represents the smallest possible architecture for our chosen range of architectures, consisting of a linear encoding, one non-linear layer followed by a linear decoding layer.
@@
For the tests shown in figure 8 of the main paper we varied the training data set size, reducing the number of time steps in the training data set to $1000$, while keeping the number of training iterations constant. The largely unchanged overall performance indicates that the original data set could be reduced in size without impeding the performance of the trained models.
@@
We computed extra- and interpolative test sets for the domain length $\mathcal{X}$. The extrapolation uses $\mathcal{X}=2\pi*[4.8,11.2]$, whilst interpolation is done for $\mathcal{X}=2\pi*[6.8,9.2]$. For our $\mathcal{L}_{2}$ results, we initialized autoregressive runs with 20 initial conditions for each $\mathcal{X}$, accumulated the $\mathcal{L}_{2}$ for sequences of $40$ steps, and averaged over all extra- and interpolative cases. The standard deviation is computed over the total set of tests. The time until decorrelation was calculated by running an autoregressive inference sequence. The steps were counted until the cross-correlation between the inferred state and the reference dropped below $0.8$. The statistics of the reached step-counts were gathered over all test cases, just like for the $\mathcal{L}_{2}$. The divergence time was defined as the number of inference steps until $\mathcal{L}_{2}>500$. Again, statistics were gathered over all test cases.
@@ 4.2 Wake Flow
Our second system is a two-dimensional flow around a cylinder, governed by the incompressible Navier-Stokes equations. The Wake Flow (WAKE) dataset consists of Karman-Vortex streets with varying Reynolds numbers, uses numerical solutions of the incompressible Navier-Stokes equations
@@
The ground truth data was simulated with operator splitting using a Chorin-projection for pressure and second-order semi-Lagrangian advection [8, 17]. The training data is generated with second-order advection, while learning tasks use a truncated spatial resolution (4×4$\times$4 ×), and first-order advection as a correction base solver. Simulations were run for a rectangular domain with a 2:1 side aspect ratio, and a cylindrical obstacle with a diameter of $L_{D}=0.1$ placed at $(0.5,0.5)$ with a constant bulk inflow velocity of $U=1$. The Reynolds number is defined as $Re=\frac{UL_{D}}{\nu}$. The physical behavior is varied with the Reynolds number by changing the viscosity of the fluid. The training data set contains 300 time steps for six Reynolds numbers $Re=[97.5,195,390,781.25,1562.5,3125]$ at resolution $[256,128]$. For learning tasks, these solutions are down-sampled by 4×4$\times$4 × to $[64,32]$. The solver for the correction uses the same solver with a more diffusive first-order advection step. The numerical solves and network training were performed in PyTorch [45].
@@
This test case employs fully-convolutional residual networks [23]. The architectures largely follow the KS setup: the neural networks have a ResNet structure, contain an additive skip connection for the output, and use a certain number of ResBlocks with a fixed number of features. The smallest network uses 1 ResBlock with 10 features, resulting in 6282 trainable parameters. The medium-sized network has 10 ResBlocks with 16 features and 66,178 parameters, while the large network has 20 blocks with 45 features resulting in more than 1 milllion (1,019,072) parameters.
@@
The networks were trained for three steps with 30000 iterations each, using a batch size of 3 and a learning rate of $1\times 10^{-4}$. The training curriculum increases the number of unrolled steps (parameter $m$ of the main text) from $m=1$, to $m=4$ and then $m=16$. Each stage applies learning rate decay with a factor of $1/10$. $\mathcal{L}_{2}$ errors are computed and accumulated over 110 steps of simulation for 12 test cases with previously unseen Reynolds numbers: three interpolative ones $Re=[2868.5,2930,2990]$ and nine extrapolative Reynolds numbers $Re=[3235,3174,3296,3845,3906,3967,4516.5,4577.5]$.
@@ 4.3 Kolmogorov Turbulence
Numerical data:
The two-dimensional KOLM system is also governed by the incompressible Navier-Stokes equations, with the addition of a forcing term $\mathbf{f}$. We use a more involved semi-implicit numerical scheme (PISO by [26]) for this setup. Similarly to the WAKE system, the KOLM learning cases are based on a spatiotemporal resolution truncation with a ratio of 4×4$\times$4 × in both space and time, and the Reynolds number is varied across training and test cases. The physical domain size was set to $L_{x}=L_{y}=2\pi$ and discretized by a $128\times 128$ grid. The timestep was set to $\Delta\tilde{t}=0.005$ for the high-resolution ground truth dataset, which maintained a Courant number smaller than $0.5$. The training dataset is based on a Reynolds number variation of $Re=[300,400,500,700,800,900]$. For each Reynolds number, $6000$ frames were added to the dataset. The numerical solves and network training were performed in Tensorflow [1].
@@
The KOLM networks are also based on ResNets. In each ResNet block, data is processed through convolutions and added to a skip connection. In contrast to KS and WAKE architectures, the number of features varies throughout the network. Four different network sizes were deployed, details of which are listed in Table 3. The network sizes in the KOLM case range from $32$ thousand to $1$ million.
@@
A spatiotemporal downsampling of the ground truth data formed the basis of the training trajectories. Thus, training operated on a 4×4$\times$4 × downsampled resolution with $4$ times larger time-steps, i.e. $32\times 32$ grid with $\Delta t=0.02$. The networks were trained with a curriculum that incrementally increased the number of unrolled steps such that $m=[1,2,4]$, with an accompanying learning rate schedule of $[10^{-4},10^{-5},10^{-6}]$. Additionally, a learning rate decay with factor $0.9$ after an epoch of $36$ thousand iterations. For each $m$, 144 thousand iterations were performed. The batch size was set to $1$.
$\mathcal{L}_{2}$ errors are computed and accumulated over 250 steps of simulation for 5 test cases with previously unseen Reynolds numbers: $Re=600$ for interpolation and $Re=1000$ for extrapolation.
Compressible Aerofoil flow (AERO):
In the final physical system, we investigate compressible flow around an aerofoil, a task centered on pure prediction. We utilize the open-source structured-grid code CFL3D [50, 49] to solve the compressible Navier-Stokes equations, thus generating our ground truth dataset. Here, we vary the Mach number while keeping the Reynolds number constant. Details of this system are found in appendix B.
Figure 3: Visualizations of our physical systems, from left to right: KS equation state over time, KOLM vorticity field, WAKE vorticity field, AERO local Mach number field visualization for different upstream Mach numbers ${}_{\infty}$
@@ 5 Results
Our results compare the three training methods ONE, NOG, and WIG in various scenarios. The figure titles mark the physical system, whilst subscripts represent the network architecture (graph: graph network, conv: convolutional ResNet), and superscripts differentiate between correction (corr) and prediction (pred) tasks. For each test, we train multiple models (typically 8 to 20) for each setup that differ only in terms of initialization (i.e. random seed). The evaluation metrics were applied independently for outputs generated by each of the models, which are displayed in terms of mean and standard deviations below. This indicates the expected performance of a training setup and the reliability of obtaining this performance. A Welch’s test for statistical significance was performed for the resulting test distributions, and p-values can be found alongside the tabulated data in E.

### v2

@@ 3.1 Kuramoto-Sivashinsky
This equation is a fourth-order chaotic PDE. It is a popular testbed for machine-learning based studies of chaotic systems. Its time-evolution was targeted by a series of previous works [97, 27, 84, 10, 71, 55, 49, 53]. The domain size, which leads to more chaotic behavior and shorter Lyapunov times for larger values [23], was varied across training and test data sets. The ground truth solver uses an exponential RK2 integrator [18]. In order to create a challenging learning scenario, the base solver for correction resorts to a first-order version that diverges within 14 steps on average. We base most of our empirical tests on this Kuramoto-Sivashinsky (KS) case since it combines challenging learning tasks with a small computational footprint. The KS equation is a fourth-order stiff PDE governed by
@@
with simulation state $u$, time $t$, and space $x$ living in the periodic domain of size $\mathcal{X}$. The dynamics exhibit a highly chaotic behavior. The domain was discretized with $N_{x}=48$ grid points such that the first computational node lies at $x_{0}=0$ and the last one at $x_{N_{x}}=\mathcal{X}*(1-1/N_{x})$. The timesteps were set to $1$. The physical domain length $\mathcal{X}$ is the critical parameter in the KS equation. The training dataset was computed for a range of $\mathcal{X}=2\pi*[5.6,6.4,7.2,8,8.8,9.6,10.4]$. To calculate the dataset, a randomized initial condition $u_{0}=cos(x)+\alpha sin(x)$, with $\alpha$ being sampled randomly from $[-0.5,0.5]$. A sequence of $5000$ steps was then recorded for each domain length. The numerical solves and network training were performed in PyTorch [64]. To calculate the dataset, a similarly randomized initial condition $u_{0}=cos(x)+\alpha sin(x)$ was used. After an initial burn-in time of 100 steps, a sequence of $5000$ steps was recorded for each domain size. The numerical solves and network training were performed in PyTorch [64]. The characteristic timescale of the KS system is based on the Lyapunov timescale of chaotic systems. This timescale $t_{s,\text{KS}}=\frac{1}{\lambda_{1}}$ requires knowledge of the largest Lyapunov exponent $\lambda_{1}$. We calculate this exponent using the Jacobian of the discretized PDE [67], which is readily available due to the differentiability of the solver. Let us denote the KS solver and its Jacobian as $u^{t+1}=\mathcal{S}_{\text{KS}}(u^{t})$, $J^{t}=\frac{\partial u^{t+1}}{\partial u^{t}}=\frac{\partial\mathcal{S}_{\text{KS}}(u^{t})}{\partial u^{t}}$. An initial perturbation $\epsilon^{0}$ is then propagated alongside the tangent of the trajectory such that
$\tilde{\epsilon}^{t+1}$ $=J^{t}\epsilon^{t},$ (19)
$\epsilon^{t+1}$ $=\frac{\tilde{\epsilon}^{t+1}}{||\tilde{\epsilon}^{t+1}||}.$ (20)
The largest Lyapunov exponent can then be calculated by estimating the exponential growth rate of the perturbation, which is achieved by averaging the normalization factor
$\lambda_{1}=\frac{1}{T}\sum_{t}^{T}\ln||\tilde{\epsilon}^{t}||,$ (21)
for a sufficiently long horizon $T$. We use this method to calculate the average Lyapunov exponents for the range $\mathcal{X}$ used in our training dataset. Setting $T=1000$ after discarding the first transient period of 100 steps for the calculations of $u_{0}$ and $\epsilon_{0}$, the average exponent evaluates to $\lambda_{1}=0.0804$. The timescale of our discretized system is thus $t_{c,\text{KS}}=12.4$, which corresponds to $\approx 12$ timesteps of the neural simulator.
@@
We use two architectures for learning tasks with the KS case: a graph-based (GCN) and a convolutional network (CNN). Similar networks were previously used on the KS equation in the case of GCNs [10, 20], while CNN based architectures were used in several other works [27, 84, 71, 55, 53]. Stachenfeld et al. [84] compared several types of convolutional networks, including ResNets and U-Nets, on the KS as well as Navier-Stokes systems. More complex convolutional architectures on the KS system were also investigated by Geneva et al. [27], who proposed an encoder-decoder based CNN, or Qu et al. [71], who split the CNN into linear and nonlinear parts for modeling of (non-)linear PDE s. In our study, both networks types follow a ResNet-like structure [33], have an additive skip connection from input to output, and are parameterized in terms of their number of features per message-passing step or convolutional layer, respectively. They receive 2 channels as input, a normalized domain size $\mathcal{X}$ and $u$, and produce a single channel, the updated $u$, as output. The GCN follows the hierarchical structure of directional Edge-Conv graph nets [101, 41] where each Edge-Conv block is comprised of two message-passing layers with an additive skip connection. The message-passing concatenates node, edge, and direction features, which are fed to an MLP that returns output node features. At a high level, the GCN can be seen as the message-passing equivalent of a ResNet. These message-passing layers in the GCN and convolutional layers in the CNN are scaled in terms of whole ResBlocks, i.e., two layers with a skip connection and leaky ReLU activation. The networks feature additional linear encoding and decoding layers for input and output. The architectures for GCN and CNN were chosen such that the parameter count matches, as listed in Table 2.
@@
For the tests shown in Figure 11 we varied the training data set size, reducing the number of time steps in the training data set to $1000$, while keeping the number of training iterations constant. The largely unchanged overall performance indicates that the original data set could be reduced in size without impeding the performance of the trained models.
@@
We computed extra- and interpolative test sets for the domain length $\mathcal{X}$. Similarly to the training dataset, the testset is based the same parameterization for the initial conditions. Again, the simulations feature a burn-in time until their dynamics are fully developed. The extrapolation uses $\mathcal{X}=2\pi*[4.8,11.2]$, whilst interpolation is done for $\mathcal{X}=2\pi*[6.8,9.2]$. For our $\mathcal{L}_{2}$ results, we selected 20 fully developed stated the test dataset and started autoregressive runs with the learned simulators on each $\mathcal{X}$. We then accumulated the relative errors for sequences of $12$ steps, corresponding to $1t_{c,\text{KS}}$. We averaged the losses over all extra- and interpolative cases and computed the standard deviation over the total set of tests.
@@ 3.2 Wake Flow
Our second system is a two-dimensional flow around a cylinder governed by the incompressible Navier-Stokes equations. This classical fluid mechanics setups also frequently used in a scientific machine learning context. See [57, 40, 93, 65, 74, 58] for some examples. Our Wake Flow (WAKE) dataset consists of Karman-Vortex streets with varying Reynolds numbers, uses numerical solutions of the incompressible Navier-Stokes equations
@@
The ground truth data was simulated with operator splitting using a Chorin-projection for pressure and second-order semi-Lagrangian advection [11, 25]. The training data is generated with second-order advection, while learning tasks use a truncated spatial resolution (4×4$\times$4 ×), and first-order advection as a correction base solver. Simulations were run for a rectangular domain with domain size $N_{x},N_{Y}=[2,1]$. A cylindrical obstacle with a diameter of $L_{D}=0.1$ placed at $(0.5,0.5)$. The streamwise boundaries use a constant bulk inflow velocity of $U=1$ at the inlet and Neumann conditions $\frac{\partial u}{\partial x}=0$ at the outlet. Inviscid, impermeable walls were used for both spanwise boundaries. The Reynolds number is defined as $Re=\frac{UL_{D}}{\nu}$. The physical behavior is varied with the Reynolds number by changing the viscosity of the fluid. The training data set contains 300 time-steps with $\Delta t=1$ for six Reynolds numbers $Re=[97.5,195,390,781.25,1562.5,3125]$ at resolution $[256,128]$. For learning tasks, these solutions are down-sampled by 4×4$\times$4 × to $[64,32]$. The solver $\mathcal{S}$ for the correction setup replaces the second-order advection with first-order semi-lagrangian advection, which introduces more numerical dissipation [85]. The numerical solves and network training were performed in PyTorch [64]. The characteristic timescale of the cylinder flow $t_{c,\text{WAKE}}$ is based on the vortex shedding frequency $f_{s}$. This frequency is calculated using the Fourier transform of a sample point downstream of the cylinder at [32,16] in the downsampled dataset. A time history of $K=500$ steps is collected, and the dominant vortex shedding frequency is extracted by
$f_{s}=\arg\max\hat{u}_{y}[f],\quad\text{where}\quad\hat{u}_{y}[f]=\sum_{k=0}^{K}u_{y}[k]e^{2\pi\iota kf/K}.$ (23)
Following this analysis, the average vortex shedding frequency in our dataset is $f_{s}=0.01$ which gives a characteristic timescale of $t_{c,\text{WAKE}}=\frac{1}{f_{s}}=100$.
@@
This test case employs fully-convolutional residual networks [33]. Our setup is an extension of the networks used in Um et al. [93] with respect to parameter count. Similar networks are a popular choice for the WAKE flow and are used in [57, 58]. Further development by Lee et al. [40] looked into multiscale CNNs on a range of physical systems including WAKE flow. CNN-based encoder-decoder architectures for similar scenarios were studied by Peng et al. [65] and Ranande et al. [74]. Our CNNs can also be seen as a two-dimensional extensions of those used the KS setup: the neural networks have a ResNet structure, contain an additive skip connection for the output, and use a certain number of ResBlocks with a fixed number of features. The smallest network uses 1 ResBlock with 10 features, resulting in 6282 trainable parameters. The medium-sized network has 10 ResBlocks with 16 features and 66,178 parameters, while the large network has 20 blocks with 45 features resulting in more than 1 million (1,019,072) parameters.
@@
The networks were trained for three steps with 30000 iterations each, using a batch size of 3 and a learning rate of $1\times 10^{-4}$. The training curriculum increases the number of unrolled steps (parameter $m$ of the main text) from $m=1$, to $m=4$ and then $m=16$. Each stage applies learning rate decay with a factor of $1/10$. Metrics are computed and accumulated over 32 steps, i.e. ca. $1/3t_{c,\text{WAKE}}$, for 8 test cases with previously unseen Reynolds numbers: three interpolative ones $Re=[2868.5,2930,2990]$ and five extrapolative Reynolds numbers $Re=[3235,3296,3906,4516.5,4577.5]$.
@@ 3.3 Kolmogorov Turbulence
Thirdly, we study a periodic two-dimensional Kolmogorov Flow (KOLM) [30] also following the incompressible Navier-Stokes equations. This setup is a widely used fluid flow case in scientific machine learning [37, 88, 82, 49, 46, 43, 4].
In contrast to the WAKE flow, KOLM introduces a forcing $\mathbf{f}$ in equation (22). The additive forcing causes the formation of a shear layer, whose instability onset develops into turbulence [30].
Numerical data:
@@
$Re=\frac{\sqrt{a_{x}}}{\nu}\Big{(}\frac{L_{y}}{2\pi}\Big{)}^{\frac{3}{2}}.$ (24)
The training dataset is based on a variation of the viscosity $\nu$, which results in a Reynolds number variation of $Re=[300,400,500,700,800,900]$. After initializing the simulations with a profile matching $\mathbf{f}*0.1$ across the entire domain, a burn-in time of $15000\Delta\tilde{t}$ was run for the high-res simulations. Afterward, $6000$ frames with $\Delta t=4\times\Delta\tilde{t}$ were added to the dataset. The numerical solves and network training were performed in Tensorflow [2]. To choose a characteristic timescale for the KOLM system that is consistent with the Reynolds number definition, we set
$t_{c,\text{KOLM}}=\sqrt{\frac{L_{y}}{2\pi*a_{x}}}.$ (25)
This definition leads to $t_{c,\text{KOLM}}=1$, which corresponds to 50 timesteps of the neural simulator since it uses a timestep of $4\times\Delta\tilde{t}$.
@@
The KOLM networks are also based on ResNets. Similar architectures were also used on a KOLM system in [37], and in [88, 43] for baseline comparison. Kochkov et al. [37] used it to study a range of unrolled systems, including one that is equivalent to our correction cases. Lippe et al. [49] used the KOLM flow as a two-dimensional testbed for their method based on convolutional U-Nets, and compared it to standard U-Nets, as well as CNNs in corrections setups. Azzizadenesheli et al. [4] based their study of neural operators on the KOLM system, and their range of networks includes operator formulations of graph networks as well as convolutional U-Nets. ResNets are structured into blocks, where data is processed through convolutions and added to a skip connection. In contrast to KS and WAKE architectures, the number of features varies throughout the network. Four different network sizes were deployed, details of which are listed in Table 3. The network sizes in the KOLM case range from $32$ thousand to $1$ million.
@@
A spatiotemporal downsampling of the ground truth data formed the basis of the training trajectories. Thus, training operated on a 4×4$\times$4 × downsampled resolution with $4$ times larger time-steps, i.e. $32\times 32$ grid with $\Delta t=0.02$. The networks were trained with a curriculum that incrementally increased the number of unrolled steps such that $m=[1,2,4]$, with an accompanying learning rate schedule of $[10^{-4},10^{-5},10^{-6}]$. Additionally, a learning rate decay with factor $0.9$ after an epoch of $36$ thousand iterations. For each $m$, 144 thousand iterations were performed. The batch size was set to $1$. $\mathcal{L}_{2}$ errors are computed and accumulated over 250 steps, corresponding to $5t_{c,\text{KOLM}}$. The procedure is repeated for 5 test cases each and previously unseen Reynolds numbers: $Re=600$ for interpolation and $Re=1000$ for extrapolation.
3.4 KS in 3D:
High-dimensional problems arise in many applications, such as in fluid mechanics in the form of LES. As a representative of higher-dimensional chaotic problems, we use a 3D extension of the KS equation. Once again, we vary the domain size in training and testing datasets. The KS equation in 3D describes a scalar-valued field $u$ in three dimensions. Similarly to the one-dimensional counterpart, it is highly chaotic. We use a cubic domain of size $\mathcal{X}\times\mathcal{X}\times\mathcal{X}$ with periodic boundaries in all directions. The equation is given by
$\frac{\partial u}{\partial t}+\big{(}\nabla u\big{)}^{2}+\nabla^{2}u+\nabla^{4}u=0.$ (26)
Numerical Data:
The ground truth solver uses an exponential RK2 integrator [18]. The domain was resolved by 32 cells in each direction. The correction is based on a first-order integrator that diverges over 14 timesteps. The networks are thus tasked to stabilize the solver and close the gap to the higher-order solution.
Network Architecture:
We focus on training correction networks for this setup, and use a ResNet [33] architecture that represent a 3D extension of the previously used 1D and 2D variants. The network features 4 ResNet blocks with 16 latent features, amounting to 55k parameters.
Network Training and Evaluation:
For the training dataset, the domain sizes were varied such that $\mathcal{X}=[12,13,15,17,19,21,22]$, and trajectories of 200 timesteps based on 5 initial conditions were run for each domain size. To construct an initial condition, we first draw a random superposition parameterized $\sin$ and $\cos$ functions, in line with the 1D case. Then, a burn-in simulation of $500$ steps was run to ensure de-correlation from these superpositions. The resulting frame is used as an initial condition for the dataset recording. Across the training dataset, we estimate the characteristic timescale based on the first Lyapunov exponent to be $t_{c,\text{KS}_{\text{3D}}}=1.8$. This is equivalent to 18 timesteps of size $\Delta t=0.1$. We trained three networks for each training variant, where each of the three used a different random seed to generate the initial network weights. All networks were trained for 50,000 steps. The unrolled versions used a horizon of $m=4$. The data-generation procedure is repeated for interpolative and extrapolative test sets, which operate on $\mathcal{X}=[14,16,18,20]$ and $\mathcal{X}=[10,11,23,24]$ respectively. The temporal horizon for testing was chosen to be 20 timesteps, which is equivalent to $1.1t_{c,\text{KS}_{\text{3D}}}$. The test losses are computed across 10 random initial conditions for each domain size.
3.5 AERO:
As a more complex physical system, we investigate compressible flow around an aerofoil, a task centered on pure prediction. We utilize the open-source structured-grid code CFL3D [78, 77] to solve the compressible Navier-Stokes equations, thus generating our ground truth dataset. Here, we vary the Mach number Ma while keeping the Reynolds number constant. Details of this system are found in C.
@@ 4 Results
Our results compare the three training methods ONE, NOG, and WIG in various scenarios. The figure titles mark the physical system, whilst subscripts represent the network architecture (graph: graph network, conv: convolutional ResNet), and superscripts differentiate between correction (corr) and prediction (pred) tasks. In the following sections, we will report on convolutional networks for the KS, WAKE, KOLM, and 3D-KS systems, graph networks on the KS system as well as U-nets in the Compressible Aerofoil flow (AERO) case. Additional U-net results for KS are presented in D.
For each test, we train multiple models (typically 8 to 20) for each setup that differ only in terms of initialization (i.e. random seed). When reporting relative errors, we use the following definition for the relative error of a single prediction
$\mathcal{L}_{\mathrm{rel}}^{t}=\frac{\mathrm{avg}\big{[}k(\mathbf{u}^{t}-\tilde{\mathbf{u}}^{t})\big{]}}{|k(\mathbf{u}^{t})|_{\infty}},\quad\text{with}\quad k(\mathbf{u}^{t})=\sum_{c}^{D}(\mathbf{u}^{t}_{c})^{2},$ (27)
where $D$ denotes the number of spatial dimensions. The $k$ operation is equivalent to calculating the kinetic energy for each grid cell and thus reduces the channel dimension of the data. The stepwise relative error is then averaged across a testing trajectory for plotting such that
$\mathcal{L}_{\mathrm{rel}}=\frac{1}{N_{T}}\sum_{t}^{N_{T}}\mathcal{L}_{\mathrm{rel}}^{t}$ (28)
The evaluation metrics were applied independently for outputs generated by each of the models, which are displayed in terms of mean and standard deviations below. This indicates the expected performance of a training setup and the reliability of obtaining this performance. A Welch’s test for statistical significance was performed for the resulting test distributions, and p-values can be found alongside the tabulated data in F.
Figure 4: Inference accuracy measured in $\mathcal{L}_{\mathrm{rel}}$ for correction setups on KS, KOLM, and WAKE systems; displayed models were trained with ONE(brown), NOG(blue), WIG(orange); across network architectures (graph networks for KS, conv-nets otherwise), and network sizes WIG has lowest errors; one 32k ONE model diverged in the KOLM case that NOG and WIG kept stable